OpenStudy (anonymous):

What is the greatest common factor of 24 and 38? 2 4 6 12

OpenStudy (anonymous):

Do you know how to factor each number?

OpenStudy (anonymous):

nope.... :/

OpenStudy (anonymous):

12

OpenStudy (anonymous):

No its not 12.

OpenStudy (anonymous):

omg yeah sorry it's not 12 nevermind I saw another number

OpenStudy (anonymous):

24= 2^3*3 38= 2*19

OpenStudy (anonymous):

its 2.

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Please explain, not just discussing among yourself...

OpenStudy (anonymous):

Lol. I did explain ?

OpenStudy (anonymous):

How did you two get 2

OpenStudy (anonymous):

thanks! this is a final exam for summer school so i need every question explained and it to be right.... i get how to do this now

OpenStudy (anonymous):

24= 2^3*3 38= 2*19

OpenStudy (anonymous):

You're welcome ! :)

OpenStudy (anonymous):

greatest common factor is the number that's COMMON BETWEEN THE NUMBERS for example this one.. 12 wasn't common in 38 so it doesn't count. even though it's common on 24.

OpenStudy (anonymous):

It's not the number that's common that's GCF, it's the product of the common factors to get GCF

OpenStudy (anonymous):

well @cydnimarie explained...

OpenStudy (anonymous):

Example 144 = 2^4 • 3^2 216 = 2^3 • 3^3 GCF = 2^3 • 3^2 = 72

OpenStudy (anonymous):

@TURITW I explained and showed work for it and the answer is 2.

OpenStudy (anonymous):

Nevermind, all matters is fine as long as kalyzza123 understands

OpenStudy (anonymous):

She said she understood.